40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 12722     Entry time: Mon Jan 16 18:54:01 2017
Author: rana 
Type: Update 
Category: SUS 
Subject: BS: whitening re-engaged 

Found that the BS whitening was off. Gautam says that "it has always been that way" and "there's nothing in the elog about this" and "I have no special relationship with Putin".

I looked at DV and DTT while turning the OSEM whitening back on. As expected, the sensor noise improved by 10x above 10 Hz. The time series shows no problems - its just less fuzzy now.

All OSEM spectra after the switch show on upper panel of plot. Lower panel shows comparison of BS UL before/after. To rotate the DTT PDF landscape output I typed this:

pdftk BS-white.pdf cat 1N output BSwhite.pdf

"if you see something, do something"

Attachment 1: BSwhite.pdf  28 kB  | Hide | Hide all
BSwhite.pdf
ELOG V3.1.3-